In a group of 270 children, there are twice as many boys as girls. How many boys are there?

Answer by laoman(51)

let b rep no of boys, g rep no of girls
from the statement
b = 2*g
2*g+g = 270
3*g = 270
g = 270/3
g = 90
b = 2* 90 = 180 BOYS
QED